OpenStudy (anonymous):

I tried working this problem out and I know someone tried to help me earlier but I just dont understand it HELP!! Rusty has a boat goes 12 mi. an hour. Find the current speed of the river if he can go 3 mi. upstream in the same time as he goes 5 mi. downstream.

OpenStudy (anonymous):

ok, distance= rate * time right?

OpenStudy (anonymous):

yes, but it asks to find the current speed...so where does that come in to play?

OpenStudy (anonymous):

I am not good with word problems at all :(

OpenStudy (anonymous):

ok, well my rate upstream is 12-x and downstream is 12+ x where x is my speed of current

OpenStudy (anonymous):

solve for time in the first equation i gave you and you have distance/rate= time

OpenStudy (anonymous):

you know that the time for going upstream 3 miles, is equal to going to downstream 5 miles

OpenStudy (anonymous):

so, 3/(12-x)=5(x+12) now just solve for x

OpenStudy (anonymous):

ok, im working on it lol ill get back to you on this one

OpenStudy (anonymous):

I got the current speed to be 12 miles per hour....is this correct?

OpenStudy (anonymous):

no,

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

i suck at this

OpenStudy (anonymous):

what did you do to solve for x?

OpenStudy (anonymous):

one second, i think i know what i did wrong

OpenStudy (anonymous):

oh wow i am totally lost :( first i multiplied both sides by 12 -x and i got 36 - 3x = 60 - 5x / x + 12 is this right so far?

OpenStudy (anonymous):

multiply both sides by the lcd of both fractions

OpenStudy (anonymous):

yes now multiply by x +12

OpenStudy (anonymous):

wait no its wrong

OpenStudy (anonymous):

when multiplying 36 - 3x by x + 12 do I use the foil method?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

its not 35-3x! because qhen you multiply x-12 to both sides, it cancels on the left side

OpenStudy (anonymous):

(x-12) * (3/12 -x)....the denominator cancels out but you still have to multiply 3 by x-12

OpenStudy (anonymous):

no! thats the whole point of multiplying by x-12 so they cancel, see when you multiply you will have 3(x-12)/(x-12) now, the x-12 cancel out and you only have left

OpenStudy (anonymous):

3

OpenStudy (anonymous):

OMG!! I hate word problems......................

OpenStudy (anonymous):

it's not the word problem, you need to practice your algebra...

OpenStudy (anonymous):

no, i always have trouble with word problems. i know how to do algebra, but you dont understand...i have been working non stop on this home work assignment for 3 days straight....i am TIRED and my brain hurts....and word problems get me so confused that i make stupid mistakes in my algebra

OpenStudy (anonymous):

yeah, i know im a tutor and i help out people, and i know how it stresses people

OpenStudy (anonymous):

ok so if it cancels out and leaves 3 then you would end up with 3 = 5 / x + 12 correct?

OpenStudy (anonymous):

yeah not multiply both sides by x+12

OpenStudy (anonymous):

3x + 36 = 5 ???

OpenStudy (anonymous):

umm multiply the x-12 to the 5 on the right side

OpenStudy (anonymous):

it does not cancelt out in the right since the denominator of the right side is x+12

OpenStudy (anonymous):

3 = 5x - 60 /x + 12

OpenStudy (anonymous):

48 miles per hour!!!!!!

OpenStudy (anonymous):

please tell me i am right

OpenStudy (anonymous):

wait lol i havent done the work yet give me a minute

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

YAY!! Thank you so much for helping me and not giving up on me!!! You are the best!!

OpenStudy (anonymous):

no problem i love math!

OpenStudy (anonymous):

i do too but i hate word problems lol

OpenStudy (anonymous):

whats the use of math if we can't apply it to the world?

OpenStudy (anonymous):

that is a very good point :)

OpenStudy (anonymous):

I have ran into a big problem with this question. When I do this equation 3 / x-12 = 5 / x+12 x turns out to equal 48. But then if you do the equation this way: 3 / 12 - x = 5 / 12+x x turns out to equal 3 So which way is correct?